OpenStudy (safyousuff97):

Task 4—Solving Rational Equations Using the equation below as a model, fill in numbers in the place of a and b to create a rational equation that has an extraneous solution. x plus a over ax = b over x Part 1. Show all work to solve for x in the equation and check the solution. Part 2. Explain how to identify the extraneous solution and what it means.

OpenStudy (eliesaab):

Is it \[ \frac {x+a}{ a x}=b \]

OpenStudy (eliesaab):

Multiply the two side by a x, to get x+a = a b x This should be easy from now on

OpenStudy (eliesaab):

Subtract abx from both sides x-a bx + a=0

OpenStudy (eliesaab):

Add -a to both sides x-abx =-a

OpenStudy (eliesaab):

You should be able to find x now

OpenStudy (eliesaab):

I will give you last step x(1-ab)=-a

OpenStudy (eliesaab):

x=?